OpenStudy (anonymous):

Which relation is a direct variation that contains the ordered pair (2, 7)? y = 4x – 1 or y=7/x or y=2/7x or y=7/2x

OpenStudy (anonymous):

y=4x-1 You just plug in your (x,y) so 7=4(2)-1 you multiply the 4(2) first which gives you 8 7=8-1 then you subtract 8-1 and get 7=7 which is true